In a factory that produces earphones, workers in the production line are each paid 300 pesos a day if they finish at most 20 units. If more than 20 units are made, they are given extra 10 pesos per unit in excess of 20. Express the amount of A earned per day by a worker as a function of the number of units X made in a day.
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

domain of units (x) is x >= 0

A = 300 (if 0 <= x <= 20)

A = 300 + 10(x-20) (if x > 20)

*(fixed)

Answer by ikleyn(52803)

.

            The solutions by   @CubeyThePenguin  is INCORRECT.

            I came to bring you the correct solution.


domain of units (x) is x >= 0

A = 300 (if 0 <= x <= 20)

A = 300 + 10*(x-20) (if x > 20)

------------

Solved.